OpenStudy (anonymous):

The solution to the radical function square root of the quantity one half x plus one equals one is x = 0

OpenStudy (anonymous):

true or false?

OpenStudy (helder_edwin):

is this \[ \large \sqrt{\frac{x}{2}+1}=1 \] correct?

OpenStudy (lgbasallote):

\[\sqrt{\frac 12 x + 1} = 0\] to solve for the zero of this square both sides to get rid of thesquare root \[(\sqrt {\frac x2 + 1})^2 = 0^2\] \[\frac x2 + 1 = 0\] now subtract 1 from both sides \[\frac x2 = -1\] now multiply 2 to both sides \[\frac x2 \times 2 = -1\times 2\] \[x = -1 \times 2\]

OpenStudy (lgbasallote):

do you think this is 0?

OpenStudy (helder_edwin):

@lgbasallote you forgot the "equals 1" of the statement

OpenStudy (lgbasallote):

where does it say it is equal 1?

OpenStudy (lgbasallote):

oh yeah

OpenStudy (lgbasallote):

lemme redo that

OpenStudy (lgbasallote):

\[\sqrt{\frac x2 + 1} = 1\] square both sides \[(\sqrt{\frac x2 + 1})^2 = (1)^2\] \[\frac x2 + 1 = 1\] subtract 1 from both sides \[\frac x2 + 1 - 1 = 1 - 1\] \[\frac x2 = 0\] multiply 2 to both sides \[\frac x2 \times 2 = 0\times 2\] \[x = 0 \times 2\] do you think this equals 0?

OpenStudy (helder_edwin):

say something @iceywheniplease7

OpenStudy (anonymous):

yes?